OpenStudy (anonymous):

does these two linear equations form a linear system and if yes why? -2x + 4y + z = 2 3x - 2/y = 0

OpenStudy (anonymous):

solve x,y,z?

OpenStudy (anonymous):

can you do that if the 2nd one is missing z variable?

OpenStudy (anonymous):

let try

OpenStudy (anonymous):

btw when will linear equations form a linear system? is it when you can solve each individual variables?

OpenStudy (anonymous):

2/y make sure u write correct

OpenStudy (anonymous):

maybe 1/2(y)

OpenStudy (anonymous):

mm no its 2 divide by y

OpenStudy (anonymous):

I never see like that,let try to see 3x - 2/y = 0 3x = 2/y x = 2/y/3 x= 2/3y Substitution x= 2/3y to first equation

OpenStudy (anonymous):

ha ha me too thats why im confused

OpenStudy (anonymous):

i finished my first year certificate IT course and working now but recently just started a computer science course and one of the modules im doing is that

OpenStudy (anonymous):

mmmmmmm so i see it doesn't really solve anything....?

OpenStudy (anonymous):

I'm not done yet

OpenStudy (anonymous):

sorry I don' t know

OpenStudy (anonymous):

its ok :P